FEL Upgrade Project Weekly Brief
June 11-15, 2001

We had a successful week of user operations providing FEL light to the nanotube collaboration and preparing for
the start of next week’s first test run of laser material studies for the JTO.

Our readers should see a review article on FELs that was written by Pat O’Shea and Henry Freund and
published in the June 8 issue of Science.

Updated report material was sent to John Albertine who is editing the report from last week’s  DoD FEL
Technology Area Working Group meeting.  Our comments included our best current estimates for an FEL
development plan for scaling FELs to 100 kW and then to the MW level.

We prepared an agenda for the upcoming June 24-25 Maritime Technical Advisory Committee which was
subsequently approved by the MTAC Chair ,VADM Al Baciocco and distributed to the committee. MTAC
advises SURA/Jefferson Lab on our Navy funded FEL program.  Given the pace of developing interests in the
FEL both by the Navy and the JTO, the advice of MTAC will be timely.

Several papers related to the FEL design and performance will be presented at next week’s Particle Accelerator
Conference in Chicago.  Dave Douglas is presenting an overview on the FEL Upgrade design (see next section
for the reference); Steve Benson is giving a paper on FEL harmonic generation; Lia Merminga is giving a paper
on high average current effects in energy recovered linacs, and Charlie Sinclair is giving a paper on our ion
implantation procedures for high voltage electrodes.  This material was presented to the FEL Upgrade review
committee at the March 27 semi-annual review.

Gwyn Williams prepared a revised pre-proposal for the joint VCU/Jefferson Lab collaboration concerned with
the re-commissioning of the Helios synchrotron for nanofabrication studies.  This was submitted to the
Commonwealth Technology Research Fund office on June 14th.

WBS 3 (Beam Physics):
Analysis of difference orbit data in support of  BBU studies continued.  Preparation for PAC2001 included the
following talk on the upgrade design:  http://www.jlab.org/~douglas/FELupgrade/talks/PAC2001.ppt

WBS 4 (Injector):
Worked on the FET system to obtain preliminary results for nitrogen implanted titanium. Continued work on gun
chamber drawings.

WBS 5 (SRF):
The cavity team tested a 7 cell cavity structure which had previously shown a degraded Q symptomatic of Q
disease.  Following a bakeout, the follow-up performance test showed improved Qs.

There is some concern that the niobium used for the recent cavity tests may be more susceptible to Q disease
because it has a noticeably smaller grain structure than previously used niobium.  A careful comparison of the
chemical and physical characteristics of our existing niobium stocks is being made.

WBS 8 (Instrumentation):
OBPM boards were completed and a daughter sample and hold board designed and built for testing.  They are
currently installed in the system and appear to be working as designed.  The "auto-lock" pico motors are being
tested today.

The beamviewer assembly area is almost ready to start production.  The clean area has been constructed and the
electricians ran power to the ventilation hood and lights, so we expect this project to begin soon.

Drawings are nearly complete for the LSS connections to the hutches in labs 3 and 6.  All labs have successfully
completed high & low level certifications.

Several additional isolation and interface module VME cards have been built.  Some for testing the new timing
circuit which is nearing completion, and some for spares to other circuits.

The website has seen some changes this week as well.  An inventory of all available VME cards has been added,
along with a search function to see the location and status of each card.  Media gallery is now part of a database
making the addition of new pictures much easier.  Anyone with questions on how to do this should contact Mike
Aston.  Also in the works is a search program for the task system that will allow easy access to tasks that have
already been completed.  This will provide a useful historical data base as well as a way to prevent "re-inventing
the wheel" when two similar tasks are performed.

An additional user/camera power supply was built. System development on the "Rabbit" board is progressing
well.  Patrick will be presenting his project at the weekly I&C meeting a week from Monday.

New drawings for the OBPM and the Absolute Value/Sample and Hold Board were given to EECAD this
week.  Drawings in EECAD:  User Shutter Control Box; IPPS Main Board Schematic and Assembly; MPS
System; Analog Differential Driver Board; Pyro-Electric OBPM Buffer/Amplifier; Picomotor Relay Chassis
Wiring Diagram; VME Timing Module Schematic, Fabrication and Assembly; Absolute Value/Sample and Hold
Board Schematic, Fabrication and Assembly.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o  The drawings of both dipoles are checked and will be transferred to the folks from DULY at the Particle
    Accelerator Conference next week.
Arc Dipoles (GY, GX, GQ)
o  Advanced Energy Systems worked on the details of the leads, bus bars, manifold, alignment cartridge mounts
    and lead cover of the Arc Bend Dipole (GX).  All these parts have to be able to work in both a left hand
    configuration as well as a right hand for two positions and the left hand configuration when the magnet is flipped over for the
    remaining two positions.
o  In magnetic modeling pursuit of flat longitudinal field, the use of a Purcell gap or the widening of the field clamp
    gap failed.
o  Comparisons between the symmetric preliminary GX and the present axisymmetric design, showed that the
    central flat field zone shifts toward the narrow end by a few cm but remains nearly identical at about 57 % of the steel
    length at the highest field, thus cross checking the TOSCA and ANSYS.
o  The IR Demo is successful at about half field where the +-1 part in 10,000 flatness extends to 65% of the
    length. AES will run a magnetic model for the upgrade at 145 MeV/c (near the field the machine will run for its
    commissioning and early running).  If the results show that we get 65% flatness, we will petition to David Douglas
    to define the operating point at this field and evaluate all our results on all the magnets at this momentum.  A
    similar process was used on the IR Demo
QX (3 inch quad)
o  The core manufacturing bids were received and evaluated.  A series of questions for the vendors to clarify
    their submissions were generated.
o  The first multifaceted cuts on the chamfers of the ends of the poles reduced n =10 pole but the N=6 pole rose.  We
    are evaluating the results
QT (Trim Quad)
o  Design work continued.
o  Robin Wines continued work on the 3D model in TOSCA.
o  The Dipole Steel order is awaiting vendor response to the flatness tolerance expected of the slabs.
o  We are working on an Interface Control Document for the task definition to be used by an engineering service
    contractor for design of the arc chambers.
o  The vendor response to the copper for all the dipoles was being evaluated.
o  Butch Dillon-Townes was able to advance the vacuum design by tabulating the vacuum pump needs of the
    Upgrade and publishing them.  A few areas need further evaluation but we now have numbers we can work

WBS 10 (Wiggler):
Four end coils are now wound for the wigglers.  A meeting was held to finalize plans for the measurement of the
wiggler.  The designer for the Hall probe carriage is on travel this week and will finish up the drawings next
week.  The 3D Hall probe is due next week.  A plan for calibrating the probe was accepted at the meeting.

WBS 11 (Optics):
The optical cavity assembly design has progressed in the following areas: we've made changes to the placement
of the relative cavity length lvdt, and added an attachment to the top of the gimbal yoke to facilitate installation an
removal.  We've increased the size of the openings across the gimbal support frame to facilitate removal of the
deformable mirrors.  K. Jordan kindly offered to supply a stepper motor that will be used for the y motion of the
assembly.  Work is transitioning to the detailing phase, with calculation of loading on the rails (used for y and z
motion), thickness of support structures, etc.

We received our Veeco laser interferometer, and scheduled installation for July 12 & 13.  We defined the
assembly of parts for the outgassing tests. (Thanks to T. Siggins & D. Bullard)

The following work was done this week to support Operations: Assisted in installation and check out of new
O-BPM detectors.  The rest of our broadband mirrors were received and we are putting them through QC.

We assisted the JTO-funded research team from NRL with initial checkout and integration of their diagnostics
with our set-up in preparation for next week’s start of the JTO run for “lethality” measurements.  We installed the
majority of our diagnostics.  We will complete the sample translation when the stage arrives on Monday.  We
ordered more material for the experiments, and made arrangements to receive other materials from NRL.

The nanotube team (CWM, NASA, Penn State) was given approximately 20 hours of run time which they used
very successfully for nanotube production.  They believe that they have generated sufficient material for extensive
purification and characterization of the fibers.

Privacy and Security Notice